OpenStudy (anonymous):

find csc 0 if tan 0= 3/4. ** the 0 has a - through it.

OpenStudy (mertsj):

Just use x instead of theta. Find cscx if tan x = 3/4

OpenStudy (mertsj):

tan x = 3/4, sin x = 3/5, csc x 5/3